OpenStudy (anonymous):

im supposed to use the table to write an equation that relates the x and y axis. the x axis is the lunch cost (L) and the y axis is the number of riders (N). the table is RIDERS (n)- 1 | 2 | 3 | 4 | 5 | 6 |7 |8 |9 Lunchcost- $4.25|8.50|12.75|17.00|21.25|25.50|29.75|34.00|28.25 sorry if its hard to read i tried my best!

OpenStudy (ranga):

Each time the number of riders goes up by 1 you will notice the lunch cost goes up by $4.25. That means it is a straight line.

OpenStudy (anonymous):

can you explain this please? what is the equation?

OpenStudy (ranga):

Lost what I just typed... Find the slope of the line: m = (8.50 - 4.25) / (2 - 1) = 4.25 Equation of the line is: y = mx + b = 4.25x + b The point (1, 4.25) lies on that line. Therefore, 4.25 = 4.25(1) + b b = 0 So the equation of the line is: y = 4.25x Here they want you to use L in the place of y and R in the place of x L = 4.25R is the equation.

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

do you have time for another one? :) the next problem says "use the equation to find the lunch cost for 25 riders"

OpenStudy (ranga):

Just plug in R = 25 in the equation L = 4.25R and calculate L.

OpenStudy (anonymous):

can you explain that? sorry.

OpenStudy (anonymous):

hello?

OpenStudy (ranga):

For part a) we found the equation of the line. It was L = 4.25 * R For part b) they are asking what is L when R = 25. Just plug in R = 25 in the equation and compute L.